OpenStudy (anonymous):

x+5/5x = 7/5 solve and check answer

OpenStudy (anonymous):

Assuming the form is \[\frac{x+5}{5x}=\frac{7}{5}\] Multiply both sides by 5x \[x+5=7x\] Subtract x from both sides and divide by 6 \[x=\frac{5}{6}\]

OpenStudy (anonymous):